# Dataset Growth

### v1

@@ 3.3 Overview of InfoGrowth
Following the theoretical analysis, we design our method InfoGrowth as shown in Fig. 2. Given an input data stream, each data point is first encoded by a backbone model (we use BLIP here) to become an embedding. Then in the embedding space, we design an efficient way to estimate the quality of the data point according to already cleaned data. Clean ones are added to the dataset with calculated "gain", while noisy ones are relabeled and re-estimated. In short, the pipeline mainly consists of a multimodal encoder model as the backbone, a cleaner for data cleaning, a gain calculator to measure the information gain, and a sampler for choosing samples. The algorithm can be summarized as Alg. 1.
@@ 4.4 Ablation
In this section, We perform extensive ablation experiments to illustrate the characteristics of InfoGrowth. If not stated otherwise, we train BLIP with 0.4M CC3M data processed with the specified ablation setting and all other settings as default, and evaluate on the COCO zero-shot retrieval task.
@@
Table 5: Neighbor k and Mean calculation. Comparison of neighborhood size and arithmetic/harmonic mean to calculate gain. k=4 with arithmetic mean performing the best.
@@
k=8 33.6 26.7 37.2 28.1
Influence of neighborhood query number k and way of calculating mean. In the gain calculator, we use the neighborhood of a given feature to calculate its information gain. Here we conduct an ablation experiment in Tab. 5 to show the influence of different k choices, TR denotes text retrieval and IR denotes image retrieval. We train the BLIP model with the dataset processed by different k values. We can see that when using the arithmetic mean, increasing k from 1 to 4 leads to a gradual performance improvement. Further increasing k to 8 shows a performance degradation. When using harmonic mean, more weight is put on closer neighbors, and increasing k gradually improves the performance, as it approximates the submodular gain on full data with higher precision.
@@
Table 7: Noise ratio. InfoGrowth provides noise resistance.
@@
Model generalization and Task generalization. As shown in Table 4.2 and  3, our processed dataset works for both BLIP(ViT) and ResNet-50 (CNN) models, with competitive performance on both multimodal and single model tasks. It suggests the effectiveness of growing a dataset within the embedding space of multimodal models is independent of model architecture and task (data) type as analyzed.
@@ 5.2 Scaling to Even Larger Scale
Figure 4: Gain decay with more data collected.
@@ 5.3 The Marginal Benefit of Data
Another interesting phenomenon of InfoGrowth is that when the total selected data amount is fixed, choosing from more candidate data doesn’t improve performance when enough data is seen (Table. 9). This implies that the data quality improvement of InfoGrowth has an upper bound because of marginal decay.

### v2

@@ 3.3 Overview of InfoGrowth
Following the theoretical analysis, we design our method InfoGrowth as shown in Fig. 2. Given an input data stream, each data point is first encoded by a backbone model (we use BLIP here) to be an embedding. Then in this embedding space, we design an efficient way to estimate the quality of the data point according to already cleaned data. Clean ones are added to the dataset with calculated "gain", while noisy ones are relabeled and re-estimated. In short, the pipeline mainly consists of: a multimodal encoder model as the backbone, a cleaner for data cleaning, a gain calculator to measure the information gain, and a sampler for choosing samples. The algorithm can be summarized as Alg. 1.
@@ 4.4 Ablation
If not stated otherwise, we train BLIP with 0.4M CC3M data processed with the specified ablation setting and all other settings as default, and evaluate on the COCO zero-shot retrieval task.
@@
Table 5: k neighbor and Mean method. Comparison of neighborhood size and arithmetic/harmonic mean to calculate gain. k=4 with mean performing the best in the experiment.
@@
k=8 33.6 26.7 33.2 27.1
Influence of neighborhood query number k and way of calculating mean. In the gain calculator, we use the neighborhood of a given feature to calculate its information gain. Here we conduct an ablation experiment in Tab. 5 to show the influence of different k choices. We train the BLIP model with the dataset processed by different k values. We can see that when using the arithmetic mean, increasing k from 1 to 4 leads to a gradual performance improvement. Further increasing k to 8 shows a performance degradation. When using harmonic mean, more weight is put on closer neighbors, and increasing k gradually improves the performance, as it approximates the submodular gain on full data with higher precision.
@@
Table 7: Noise ratio. Our method provides noise resistance.
@@
Model generalization and Task generalization. In Table 4.2, we test the performance of InfoGrowth on CLIP architecture. It suggests the effectiveness of growing a dataset within the embedding space of multimodal models is independent to model architecture.
Table 8: Video classification on MSRVTT and Image classification on ImageNet-1K for BLIP.
Dataset Scale MSRVTT Imagenet
[…]
Rand-cc3m 0.67M 23.3 42.8 53.3 57.3
cc3m 2.82M 26.0 46.3 58.0 62.5
Ours 0.67M 26.4 47.3 58.4 62.1
Ours 1.35M 28.8 49.4 60.0 62.9
In Table 8, we can see that on MSRVTT, which is video classification with different data distribution to BLIP training data, our trained BLIP on 0.67M CC3M data processed by InfoGrowth shows improved ability than trained on the original CC3M. InfoGrowth also improves over the original CC3M on ImageNet-1k zero-shot classification with 1.35M processed data. It suggests that task generalizability is inherited by InfoGrowth processed dataset.
@@ 5.3 The Marginal Benefit of Data
Figure 4: Gain decay with more data collected.
@@
Another interesting phenomenon of InfoGrowth is that when the total selected data amount is fixed, choosing from more candidate data doesn’t improve performance when enough data is seen (Table. 10). This implies that the data quality improvement of InfoGrowth has an upper bound because marginal decay.
@@ 6 Conclusion
Acknowledgements
This work is supported by Alibaba Group through Alibaba Innovative Research Program. This work is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-PhD-2021-08-008). Yang You’s research group is being sponsored by NUS startup grant (Presidential Young Professorship), Singapore MOE Tier-1 grant, ByteDance grant, ARCTIC grant, SMI grant (WBS number: A-8001104-00-00), Alibaba grant, and Google grant for TPU usage. This work is partially supported by the National Natural Science Foundation of China (62176165), the Stable Support Projects for Shenzhen Higher Education Institutions (20220718110918001), the Natural Science Foundation of Top Talent of SZTU(GDRC202131). This work was supported in part by the National Science and Technology Major Project (2021ZD0110901). We thank Haonan Wang for valuable discussions and feedback.